eBook In alphabetical order, the six most common last names in the United States are Brown, Johnson, Jones, Smith, Miller, and Williams (The World Almanac, 2012). Assume that a sample of 50 individuals with one of these last names provided the following data. Click on the datafile logo to reference the data DATA file Brown Williams Williams Williams Brown Smith Jones SmithJohnson Smith Miller Sth Brown Williams Johnson Johnson Smith Smith ohnson Brown Williams Miller ohson Williams Johnson Williams Johnson Jones SmithBrown Johnson Smith Smith Bown Jones Jones Miller Jones Jones Johnson Brown Johnson Miller Jones Smith Sith Miller Wilams Miller Smith Summarize the data by constructing the following items. a. Fll In the frequency and the relatlve frequency (to 2 declmals), and percent frequency values below NameFrequency Relative Percent Frequency frequency Brown Johnson Jones Miller Smith Williams (Total)b. Which of the following three bar graphs accurately represents the data? Common U.S. Last Names Frequency 20 15 Smith Johnsorn 10 William:s Brown Jones Miller -52. Common U.S. Last Names Frequency 20 15 Williams Brown 10 Smith Johnson MillerI Jones -53. Common U.S. Last Names Frequency 20 15 Johnson Smith 10 Brown Miller Williams Jones -5 - Select your answerc. Select the correct pie chart from the following 1. Common U.S. Last Names Johnson 20% Jones 14% Brown 14% Miller 1296 Williams 16% Smith 24% 2. Common U.S. Last Names ohnson 14 Jones 12% Brown 20% Miller 14% Williams 24% Smith 16963. Common U.S. Last Names Johnson 24% Brown 16% Jones 12% Williams 14% Miller 14% Smith 20% - Select your answer - d. Based on these data, what are the three most common last names? Select your answer

Those pictures are for one question, please answer all the questions.
Answer #1

We fill the frequency values as follows:

Names Frequency Relative Frequency Percentage frequency
Brown 7 7/50=0.14 14
Johnson 10 10/50=0.2 20
Jones 7 7/50=0.14 14
Miller 6 6/50=0.12 12
Smith 12 12/50=0.24 24
Williams 8 8/50=0.16 16
Total 50 1 100

Smith has the highest frequency of 12 and percentage of 24%. Thus histogram in option (1) is correct and pie chart in option (1) is correct.

The three common last names are Smith (freq 12), Johnson (freq 10) and Williams (freq 8).